OpenStudy (anonymous):

The cylinder shown below has a radius of 3 inches and a height of 23 inches. What is the approximate lateral area of the cylinder? Use for pi. 433.71 in2 490.29 in2 462.00 in2 216.86 in2

OpenStudy (jack1):

by lateral did you mean "outside"? sorry, i cant see the picture

OpenStudy (jack1):

If you mean the surface area, the answer is: 2x Area of Base + area of rectangle =2x (pi x r^2) + (2 x pi x r x ht)

OpenStudy (anonymous):

HERE LATERAL SURFACE AREA IS CURVED SURFACE AREA = 2*PI*R*H = 2 *3.14*3*23 = 433 INCHES (APPROX)

OpenStudy (anonymous):

thanks so much !!!

OpenStudy (anonymous):

i have another question can you please help?

OpenStudy (anonymous):

A cone with a radius of 6 feet and a slant height of 13 feet is shown below. What is the surface area of this cone? Use for pi. 942.86 square feet 113.14 square feet 245.14 square feet 358.29 square feet